OpenStudy (anonymous):

Evaluate the following and write the answer in scientific notation , (6*10^-5)(5*10^-8) @-------College Al…

myininaya (myininaya):

\[6 \cdot 5 \cdot 10^{-5+(-8)}=30 \cdot 10^{-13}\]